OpenStudy (anonymous):

What are the vertical asymptotes if there are any? f(x)= (x-4)/ x(x-4)

OpenStudy (anonymous):

\[f(x)=\frac{x-4}{x(x-4)}=\frac{1}{x}\] if \(x\neq 4\)

OpenStudy (anonymous):

only vertical asymptote is \(x=0\)